# Scaling Laws for Moral Machine Judgment in Large Language Models

### v1

@@ Abstract
Autonomous systems increasingly require moral judgment capabilities, yet whether these capabilities scale predictably with model size remains unexplored. We systematically evaluate 75 large language model configurations (0.27B–1000B parameters) using the Moral Machine framework, measuring alignment with human preferences in life-death dilemmas. We observe a consistent power-law relationship with distance from human preferences () decreasing as (, ) where is model size. Mixed-effects models confirm this relationship persists after controlling for model family and reasoning capabilities. Extended reasoning models show additional 16% improvement beyond scale effects. The relationship holds across diverse architectures, while variance decreases at larger scales, indicating systematic emergence of more reliable moral judgment with computational scale. These findings extend scaling law research to value-based judgments and provide empirical foundations for artificial intelligence governance.
@@ 1 Introduction
Recent work has established scaling laws across diverse cognitive capabilities. Model performance improves predictably with increased parameters in language modeling and understanding [16, 13]. Many sophisticated capabilities emerge spontaneously through scale alone, without explicit training [30]. These observations raise the question of whether moral judgment capabilities similarly scale with model size, following patterns analogous to other cognitive domains.
@@ 2.4 Statistical Analysis
To assess whether the observed scaling relationship reflects effects of model size rather than correlated confounding factors, we employed linear mixed-effects models with model family (DeepSeek, Llama, Gemma, Qwen, Other) as random effects, including family-specific intercepts and slopes for model size. We included two additional predictors: (1) release date, converted to a continuous variable (years) to capture temporal trends and serve as a proxy for potential improvements in training data quality and methods over time; (2) reasoning capability, coded as a binary variable indicating whether models employ extended reasoning processes (reasoning-focused models and thinking-mode variants = 1, standard models = 0). We compared nested models using likelihood ratio tests and information criteria (AIC, BIC): (A) model size only, (B) model size plus release date, (C) model size plus reasoning capability, and (D) model size, reasoning capability, and their interaction.
@@ 3.2 Robustness to Confounding Factors
To assess whether the observed scaling relationship reflects genuine effects of model size rather than confounding factors, we fitted linear mixed-effects models with model family (DeepSeek, Llama, Gemma, Qwen, Other) as random effects. This approach accounts for family-specific differences in architecture and training approaches while estimating the effect of size. Model size correlates with several factors that could contribute to improved alignment, including computational budget, architectural sophistication, and training data quality. While the first two can be partially controlled through family-specific random effects (as models within families share similar architectures and resource access), training data quality is not directly observable. We used release date, converted to a continuous variable, as a proxy for potential improvements in training data curation and quality over time. We compared nested models using likelihood ratio tests: (A) model size only, (B) model size plus release date, (C) model size plus reasoning capability, and (D) model size, reasoning capability, and their interaction (Table 1).
@@
| C | Size + Reasoning | 7.24 | 0.007 | ||
| D | Size Reasoning | 0.35 | 0.55 | ||
| and -values from likelihood ratio tests against Model A | |||||
Model comparison revealed that the best-fitting model included both model size and reasoning capability as fixed effects (Model C: AIC , BIC ; Table 1), significantly outperforming the size-only model (Model A: , , ). Critically, adding release date did not improve fit (Model B vs. A: ), suggesting that temporal improvements in training data quality do not substantially contribute to alignment beyond the effects captured by model size and family. The size reasoning interaction was also non-significant (Model D vs. C: ).
The final model (Model C; Table S3) confirmed the power-law scaling relationship while controlling for family differences. Model size showed a significant negative effect (, SE , ), corresponding to a power-law exponent consistent with the overall analysis. Family-specific patterns (Figure S1) show predominantly negative relationships between size and alignment, though statistical power for detecting these relationships within individual families varies due to differences in sample size (8–19 models per family) and parameter range coverage. Random effects captured modest family-level variation (SD, SD; Table S3), with the mixed-effects approach providing robust estimates by pooling information across families while accounting for family-specific heterogeneity.
Extended reasoning models showed significantly better alignment (, SE , ; Table S3), reducing distance by approximately 16% (95% CI: 4–26%) beyond the effect of size alone (Figure S2). This effect persists after controlling for model size, indicating that architectural or training innovations enabling extended reasoning contribute to moral alignment independently of computational scale. Comparison of coefficient estimates across all nested models is provided in Table S4.
Analysis of residuals after controlling for size showed no correlation with release date (Spearman , ; Figure S3; Table S4), further supporting that temporal improvements do not substantially contribute to alignment beyond scale and reasoning capabilities. This null result should be interpreted cautiously given the limited temporal range of our dataset (95% of models released in 2024–2025), which may constrain our ability to detect gradual improvements in training data quality over time.
@@ 4 Discussion
Our findings establish that computational scale systematically predicts moral judgment capabilities in LLMs. Across 75 model configurations spanning three orders of magnitude (0.27B–1000B parameters), we observed a consistent power-law relationship (, , ), where distance from human moral preferences decreases predictably with model size. This relationship proved robust to potential confounding factors: mixed-effects models controlling for model family confirmed that the scaling pattern persists across diverse architectures and training approaches (, ). Family-specific analyses revealed predominantly negative relationships between size and alignment across all major model families (DeepSeek, Llama, Gemma, Qwen), demonstrating that the power-law pattern is not driven by any single architectural approach or training methodology.
@@
While computational scale represents the primary determinant of moral alignment, additional factors contribute independently. Extended reasoning models demonstrated significantly better alignment (, ), reducing distance by approximately 16% beyond the effect of size alone. This indicates that architectural or training innovations enabling extended reasoning (e.g., chain-of-thought processing [31, 29] and iterative refinement [19]) enhance moral judgment independently of parameter count. In contrast, temporal improvements in training data quality, proxied by release date, did not contribute substantially to alignment after controlling for size and reasoning capability (). This null result suggests that while model size and reasoning architecture represent primary pathways to improved moral alignment, gradual refinements in data curation practices do not provide comparable benefits, at least within the temporal range of our dataset (2023–2025).
@@
The modest power-law exponent () reveals an important characteristic of moral judgment capabilities: they scale slowly with model size. A tenfold increase in parameters yields only approximately 21% improvement in alignment, indicating that moral judgment represents a particularly challenging emergent capability requiring substantial computational scale to approach human-level performance. This gradual scaling contrasts with steeper improvements observed in some other domains and, while subject to the interpretation of our distance-based metric, suggests fundamental computational constraints on acquiring human-like moral intuitions. For practical deployment, this slow scaling implies that achieving substantial improvements in moral alignment requires order-of-magnitude increases in model size or complementary approaches such as extended reasoning architectures, which provide comparable benefits (approximately 16% improvement) without proportional increases in parameter count. Beyond improved mean alignment, larger models also exhibit reduced variance in performance, clustering more tightly around the power-law trend and indicating systematic emergence of more reliable and consistent moral judgment capabilities with scale. This dual benefit of improved alignment and increased reliability suggests that computational scale enhances both the quality and predictability of moral reasoning in safety-critical applications.
@@
These findings provide quantitative foundations for risk-based deployment of LLMs in contexts requiring moral judgment. The power-law relationship enables prediction of alignment quality from model size, allowing system designers to make informed tradeoffs between computational costs and ethical reliability. For safety-critical applications where ethical failures carry substantial consequences, our results indicate that larger models or those enhanced with extended reasoning capabilities offer improved reliability. The 16% improvement provided by extended reasoning represents a complementary pathway to enhanced alignment that does not require proportional increases in parameter count. Beyond mean alignment, the observed reduction in variance at larger scales implies more predictable performance, reducing the risk of outlier judgments that could undermine trust or violate ethical standards in high-stakes contexts [26, 25, 14].
@@
Model size correlates with multiple factors including training data quality, computational budget, and architectural sophistication. While our mixed-effects analyses control for model family and examine temporal trends, fully isolating the causal effect of scale from all potential confounders requires controlled experiments that are impractical at the scale of contemporary LLMs. Training data for most models is not publicly disclosed, making it impossible to assess whether Moral Machine scenarios appeared in pretraining corpora, though the consistency of scaling patterns across models from different organizations suggests systematic relationships rather than memorization artifacts.
@@ Acknowledgments
| Intercept | 0.006 | 0.106 | 30.1 | 0.061 | 0.952 |
| Model Size (log10) | 0.012 | 2.1 | 0.013 | ||
| Release Date | 0.034 | 0.023 | 35.6 | 1.506 | 0.141 |
| Extended Reasoning | 0.027 | 68.3 | 0.008 | ||
@@
| Family: Intercept = 0.005 | |||||
| Family: Slope (Model Size) = 0.005 | |||||
@@
| Marginal (fixed effects): 0.50; Conditional (full model): 0.52 | |||||
| AIC = (REML); Log-likelihood = 68.1 | |||||
@@
| Family: Intercept | 0.042 | 0.040 | 0.005 | 0.015 |
| Family: Slope (Size) | 0.013 | 0.009 | 0.005 | 0.015 |
| Residual | 0.096 | 0.095 | 0.094 | 0.094 |
@@
| Log-likelihood | 66.8 | 67.8 | 68.1 | 65.5 |
| AIC (ML)a | ||||
| aMaximum likelihood estimates used for model comparison (Table 1) | ||||

### v2

@@ Abstract
Autonomous systems increasingly require moral judgment capabilities, yet whether these capabilities scale predictably with model size remains unexplored. We systematically evaluate 75 large language model configurations (0.27B–1000B parameters) using the Moral Machine framework, measuring alignment with human preferences in life-death dilemmas. We observe a consistent power-law relationship with distance from human preferences () decreasing as (, ) where is model size. Mixed-effects models confirm this relationship persists after controlling for model family and reasoning capabilities. Extended reasoning models show significantly better alignment, with this effect being more pronounced in smaller models (sizereasoning interaction: ). The relationship holds across diverse architectures, while variance decreases at larger scales, indicating systematic emergence of more reliable moral judgment with computational scale. These findings extend scaling law research to value-based judgments and provide empirical foundations for artificial intelligence governance.
@@ 1 Introduction
Recent work has established scaling laws across diverse cognitive capabilities. Model performance improves with increased parameters in language modeling and understanding [16, 13]. Many sophisticated capabilities emerge spontaneously through scale alone, without explicit training [30]. These observations raise the question of whether moral judgment capabilities similarly scale with model size, following patterns analogous to other cognitive domains.
@@ 2.4 Statistical Analysis
To assess whether the observed scaling relationship reflects effects of model size rather than correlated confounding factors, we employed linear mixed-effects models with model family (DeepSeek, Llama, Gemma, Qwen, Other) as random effects, including family-specific intercepts and slopes for model size. We included two additional predictors: (1) release date, converted to a continuous variable (years) to capture temporal trends and control for potential scale-independent improvements in model capabilities, such as advances in training methodology, architecture design, or alignment procedures; (2) reasoning capability, coded as a binary variable indicating whether models employ extended reasoning processes (reasoning-focused models and thinking-mode variants = 1, standard models = 0). We compared nested models using likelihood ratio tests and information criteria (AIC, BIC): (A) model size only, (B) model size and release date, (C) model size, release date, and reasoning capability, and (D) model size, release date, reasoning capability, and sizereasoning interaction.
@@ 3.2 Robustness to Confounding Factors
To assess whether the observed scaling relationship reflects genuine effects of model size rather than confounding factors, we fitted linear mixed-effects models with model family (DeepSeek, Llama, Gemma, Qwen, Other) as random effects. This approach accounts for family-specific differences in architecture and training approaches while estimating the effect of size. Model size correlates with several factors that could contribute to improved alignment, including computational budget, architectural sophistication, and advances in training methodology. While the first two can be partially controlled through family-specific random effects (as models within families share similar architectures and resource access), temporal improvements in model development are not directly observable. We used release date, converted to a continuous variable, as a proxy for potential scale-independent improvements over time, which could include advances in training practices, data curation, or alignment techniques. We compared nested models using likelihood ratio tests: (A) model size only, (B) model size and release date, (C) model size, release date, and reasoning capability, and (D) model size, release date, reasoning capability, and sizereasoning interaction (Table 1).
@@
| C | Size + Release date + Reasoning | 7.24 | 0.007 | ||
| D | Size + Release date + Reasoning + SizeReasoning | 5.05 | 0.025 | ||
| and -values from likelihood ratio tests against previous model | |||||
Model comparison revealed that adding release date did not improve fit (Model B vs. A: , ), suggesting that temporal factors in model development, which could include advances in training methodology, architectural innovations, or alignment procedures, do not substantially contribute to moral alignment beyond the effects captured by model size and family. However, adding reasoning capability significantly improved the model (Model C vs. B: , ). The full model including the sizereasoning interaction provided further significant improvement (Model D vs. C: , ).
The final model (Model D; Table S3) confirmed the power-law scaling relationship while controlling for family differences and temporal trends. Model size showed a significant negative effect (, SE , ), corresponding to a power-law exponent consistent with the overall analysis. Extended reasoning models showed significantly better alignment (, SE , ). Family-specific patterns (Figure S1) show predominantly negative relationships between size and alignment, though statistical power for detecting these relationships within individual families varies due to differences in sample size (8–19 models per family) and parameter range coverage. Random effects captured modest family-level variation (SD, SD; Table S3), with the mixed-effects approach providing robust estimates by pooling information across families while accounting for family-specific heterogeneity. Critically, the sizereasoning interaction was positive and significant (, SE , ), indicating that the advantage of extended reasoning capabilities is more pronounced in smaller models (Figure S2). This suggests that architectural innovations in reasoning provide complementary pathways to improved alignment, with particularly strong benefits when computational scale is limited, while very large models may already capture much of this reasoning capability through scale alone. Comparison of coefficient estimates across all nested models is provided in Table S4.
Analysis of residuals after controlling for size showed no correlation with release date (Spearman , ; Figure S3; Table S4) further supporting that scale-independent temporal improvements in model development do not substantially contribute to alignment beyond scale and reasoning capabilities. This null result should be interpreted cautiously given the limited temporal range of our dataset (95% of models released in 2024–2025), which may constrain our ability to detect gradual improvements in training data quality over time.
@@ 4 Discussion
Our findings establish that computational scale systematically predicts moral judgment capabilities in LLMs. Across 75 model configurations spanning three orders of magnitude (0.27B–1000B parameters), we observed a consistent power-law relationship (, , ), where distance from human moral preferences decreases predictably with model size. This relationship proved robust to potential confounding factors: mixed-effects models controlling for model family confirmed that the scaling pattern persists across diverse architectures and training approaches. Family-specific analyses revealed predominantly negative relationships between size and alignment across all major model families (DeepSeek, Llama, Gemma, Qwen), demonstrating that the power-law pattern is not driven by any single architectural approach or training methodology.
@@
While computational scale represents the primary determinant of moral alignment, additional factors contribute independently. Extended reasoning models demonstrated significantly better alignment (, ). Critically, this effect varies with model size: the sizereasoning interaction (, ) indicates that extended reasoning provides greater relative benefits in smaller models, while very large models already capture much of this capability through scale alone. This indicates that architectural or training innovations enabling extended reasoning, such as chain-of-thought processing [31, 29] and iterative refinement [19], enhance moral judgment, with this enhancement being particularly pronounced when computational scale is limited. In contrast, temporal factors in model development, proxied by release date, did not contribute substantially to alignment after controlling for size and reasoning capability (). This null result suggests that while model size and reasoning architecture represent primary pathways to improved moral alignment, scale-independent improvements over time—which could include advances in training practices, data curation, or alignment techniques—do not provide comparable benefits, at least within the temporal range of our dataset (2023–2025).
@@
The modest power-law exponent () reveals an important characteristic of moral judgment capabilities: they scale slowly with model size. A tenfold increase in parameters reduces distance from human preferences by only approximately 21%, indicating that moral judgment represents a particularly challenging emergent capability requiring substantial computational scale to approach human-level performance. This gradual scaling contrasts with steeper improvements observed in some other domains and, while subject to the interpretation of our distance-based metric, suggests fundamental computational constraints on acquiring human-like moral intuitions. For practical deployment, this slow scaling implies that achieving substantial improvements in moral alignment requires order-of-magnitude increases in model size or complementary approaches such as extended reasoning architectures. Extended reasoning provides particularly strong benefits in resource-constrained settings where deploying very large models is impractical, offering an alternative pathway to improved alignment without proportional increases in parameter count. Beyond improved mean alignment, larger models also exhibit reduced variance in performance, clustering more tightly around the power-law trend and indicating systematic emergence of more reliable and consistent moral judgment capabilities with scale. This dual benefit of improved alignment and increased reliability suggests that computational scale enhances both the quality and predictability of moral reasoning in safety-critical applications.
@@
These findings provide quantitative foundations for risk-based deployment of LLMs in contexts requiring moral judgment. The power-law relationship enables prediction of alignment quality from model size, allowing system designers to make informed tradeoffs between computational costs and ethical reliability. For safety-critical applications where ethical failures carry substantial consequences, our results indicate that larger models or those enhanced with extended reasoning capabilities offer improved reliability. Extended reasoning represents a complementary pathway to enhanced alignment that does not require proportional increases in parameter count, with particularly strong benefits when computational scale is limited. Beyond mean alignment, the observed reduction in variance at larger scales implies more predictable performance, reducing the risk of outlier judgments that could undermine trust or violate ethical standards in high-stakes contexts [26, 25, 14].
@@
Model size correlates with multiple factors including computational budget, architectural sophistication, and advances in training methodology. While our mixed-effects analyses control for model family and temporal trends in model development, fully isolating the causal effect of scale from all correlated factors requires controlled experiments that are impractical at the scale of contemporary LLMs. Training data for most models is not publicly disclosed, making it impossible to assess whether Moral Machine scenarios appeared in pretraining corpora, though the consistency of scaling patterns across models from different organizations suggests systematic relationships rather than memorization artifacts.
@@ Acknowledgments
| Intercept | 0.033 | 0.107 | 56.6 | 0.312 | 0.756 |
| Model Size (log10) | 0.016 | 5.9 | 0.001 | ||
| Release Date | 0.033 | 0.023 | 58.4 | 1.458 | 0.150 |
| Reasoning | 0.044 | 40.3 | 0.001 | ||
| SizeReasoning | 0.057 | 0.025 | 56.3 | 2.313 | 0.024 |
@@
| Family: Intercept = 0.019 | |||||
| Family: Slope (Model Size) = 0.017 | |||||
@@
| AIC = (REML) | |||||
@@
| Family: Intercept | 0.042 | 0.040 | 0.005 | 0.019 |
| Family: Slope (Size) | 0.014 | 0.009 | 0.005 | 0.017 |
| Residual | 0.095 | 0.095 | 0.094 | 0.091 |